# Products of two orthogonal projections

### v1

@@ Abstract.
We study operators that are products of two orthogonal projections. Our results complement the classical results of Crimmins and von Neumann. Particular emphasis has been given to projections associated with inner functions defined on the polydisc.
@@ 1. Introduction
for some projections $P_{1}$ and $P_{2}$ in $\mathcal{B}(\mathcal{H})$. Operators that admit the above factorizations have been examined in multiple contexts, including von Neumann’s formula for iterated products of projections and Crimmins’ analysis of products of two projections. Aronszajn, Browder, Dixnier, Kakutani, and Weiner are among the many more eminent names who have contributed to illuminate this subject (see [15, 23] for a thorough historical description). Let us revisit two characterizations from the list provided by Crimmins [5, 23] : An operator $T\in\mathcal{B}(\mathcal{H})$ is the product of two projections if and only if
@@
In this paper, given a closed subspace $\mathcal{S}$ of a Hilbert space $\mathcal{H}$, we denote the orthogonal projection of $\mathcal{H}$ onto $\mathcal{S}$ by $P_{\mathcal{S}}$. The preceding factorization of $T$ is referred to as the canonical factorization and will be crucial in the subsequent discussion. We present somewhat independent proofs of the above equivalence properties. Additionally, we offer the following new characterization: $T\in\mathcal{B}(\mathcal{H})$ is a product of two projections if and only if
@@
Suppose $T\in\mathcal{B}(\mathcal{H})$ is a product of projections. Then $T$ is necessarily a contraction, that is, $\|Th\|\leq\|h\|$ for all $h\in\mathcal{H}$. Consequently, the theory of contractions [18]— an influential and impactful theory—applies to operators that are products of two projections. This paper presents some new perspectives on the structure of contractions that are products of two projections. We begin with the celebrated canonical decompositions of contractions.
Let $T\in\mathcal{B}(\mathcal{H})$ be a contraction. The canonical decomposition of $T$ yields the orthogonal decomposition of closed subspaces $\mathcal{H}=\mathcal{H}_{u}\oplus\mathcal{H}_{cnu}$, where $\mathcal{H}_{u}$ and $\mathcal{H}_{cnu}$ reduce $T$, and $T|_{\mathcal{H}_{u}}$ is unitary and $T|_{\mathcal{H}_{cnu}}$ is cnu (that is, $T|_{\mathcal{H}_{cnu}}$ on $\mathcal{H}_{cnu}$ does not have unitary summand). Moreover, the unitary summand $\mathcal{H}_{u}$ is given by
@@
respectively, where $\bigvee$ denotes the span closure of subspaces. The above results complement the classical von Neumann’s alternating orthogonal projection formula. In fact, if $P_{1}$ and $P_{2}$ are projections, then the alternating orthogonal projection formula tells us that
@@
where $\mathcal{H}_{u}$ is the unitary part of the canonical decomposition of the contraction $P_{1}P_{2}$. In the context of canonocial decomposition, we additionally include that (see Theorem 4.2)
@@
This amounts to saying the following asymptotic properties of the cnu part of $T$:
Now we look into a more concrete infinite dimensional Hilbert space, $H^{2}(\mathbb{D}^{n})$, the Hardy space of square summable analytic functions defined on the polydisc $\mathbb{D}^{n}$. The Banach algebra of bounded analytic functions on $\mathbb{D}^{n}$ is denoted by $H^{\infty}(\mathbb{D}^{n})$. Therefore, we have
@@
for all $f\in H^{2}(\mathbb{D}^{n})$. One knows that $\|T_{\varphi}\|=\|\varphi\|_{\infty}$ for all $\varphi\in H^{\infty}(\mathbb{D}^{n})$ [9]. The function $\varphi\in H^{\infty}(\mathbb{D}^{n})$ is said to be inner if $|\varphi|=1$ a.e. on $\mathbb{T}^{n}$ in the sense of radial limits. It is known that $\varphi$ is inner if and only if $T_{\varphi}$ is an isometry on $H^{2}(\mathbb{D}^{n})$. In this case, $T_{\varphi}T_{\varphi}^{*}$ is an orthogonal projection onto $\varphi H^{2}(\mathbb{D}^{n})$. In other words, we have
@@
$P=P_{\varphi H^{2}(\mathbb{D}^{n})}.$
Given a bounded linear operator $T$ on $H^{2}(\mathbb{D}^{n})$, define the set (which appears to be a nonemtpy set)
@@
Under this notation, we also have the following:
@@
Theorem 5.3 yields an analytic characterization of products of two inner projections: Let $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ be a nonzero operator. Then $T$ is a product of two inner projections if and only if
@@
We note that, unlike $\mathcal{I}_{T}$, the set $\mathcal{J}_{T}$ has the potential to remain empty. Therefore, part of the above result also guarantees that the set $\mathcal{J}_{T}$ remains nonempty. In the final section of this paper, we illustrate some of our results through concrete examples as above.
The remaining part of the paper is arranged as follows: Section 2 primarily focuses on proving some of the well-known results, including those by Crimmins and Sebestyén. However, the more general framework we’ve provided here also emphasizes the complexity of the larger problem. In Section 3, we explain the classical concept of canonical decompositions of contractions and relate them to our theory of the product of two projections. This provides a precise description of the orthogonal decompositions applicable to the operators under consideration. Section 4 establishes an additional connection between our theory and a classical result, providing new insights into von Neumann’s alternating orthogonal projection formula. In Section 5, we examine our underlying problem from a specific standpoint. We apply our theory to projections based on inner functions defined on the Hardy space over the polydisc. We continue to work on concrete situations in Section 6, where the projections are considered to be model projections. The final section, Section 7, illustrates some of our results on products of two projections using Blaschke products.
2. Crimmins’ perspectives
In this section, we prove some results on operators that can be represented as products of two projections. As stated in the very beginning, this is a problem that has been triggered by many. We take Crimmins’ point of view and dedicate a part of the section to reproving his results. However, we address this within a marginally wider framework, anticipating more realistic uses in the structure of operators that can be expressed as the product of operators.
@@ 2. Crimmins’ perspectives
If $T=PX$ for some projection $P\in\mathcal{B}(\mathcal{H})$ and linear operator $X\in\mathcal{B}(\mathcal{H})$, then
@@
Now, $T=P_{\overline{\text{ran}T}}T$, and hence $T=P_{\overline{\text{ran}T}}PX$. This implies
$T=P_{\overline{\text{ran}T}}X,$
which completes the proof of the lemma.
@@
If $T=P_{1}XP_{2}$ for some projections $P_{1},P_{2}\in\mathcal{B}(\mathcal{H})$ and linear operator $X\in\mathcal{B}(\mathcal{H})$, then
@@
We are now presenting a result of Sebestyén [17, Corollary 2]. The present proof, compared to [17], is more elementary (the sufficient part) and is also presented for completeness. The result is clearly in line with Douglas’ range inclusion theorem.
@@
Crimmins once again receives credit for the equivalence of (1) and (4) [5, page 1595]. The present proof also is once again different. In this context, we also refer the reader to [5, Theorem 3.1]. The same paper contains results (cf. [5, Corollary 3.8]) related to parametrizations and the uniqueness of pairs of projections whose products yield a given operator.
@@ 3. Nagy-Foias and Langer’s perspective
Note that if a linear operator acting on a Hilbert space is a product of two projections, then it is necessarily a contraction. Consequently, the theory of contractions, widely recognized as being extremely rich, applies in this situation. This is a vast domain, and a thorough application of the theory of contractions to this particular class of operators and vice versa may not be entirely transparent. This requires a more in-depth investigation, which ought to be carried out. However, in this section, we will address a fundamental characteristic of the product of projections via the structure of contractions. To do that, we revisit the classic Sz.-Nagy, Foias, and Langer orthogonal decompositions of contractions [12, 19].
Let $T\in\mathcal{B}(\mathcal{H})$ be a contraction. Then there is a unique orthogonal decomposition $\mathcal{H}=\mathcal{H}_{u}\oplus\mathcal{H}_{cnu}$ such that both $\mathcal{H}_{u}$ and $\mathcal{H}_{cnu}$ reduce $T$, and $T|_{\mathcal{H}_{u}}$ is unitary whereas $T|_{\mathcal{H}_{cnu}}$ is completely non-unitary (cnu in short). This amounts to saying that $T|_{\mathcal{H}_{cnu}}$ does not have unitary summand. As a result, we get the diagonal decomposition:
@@
on $\mathcal{H}=\mathcal{H}_{u}\oplus\mathcal{H}_{cnu}$. Moreover, we have the representation of the unitary part $\mathcal{H}_{u}$ as [18, Chapt. 1, Theorem 3.2]
@@
This decomposition is popularly known as the Sz. Nagy-Foiaş-Langer or the canonical decomposition of contractions and stands as one of the most basic structures of contractions. In the following, we connect this with our theory of the product of two projections. For simplicity of notation, given projections $P_{1}$ and $P_{2}$, define the closed subspace $\mathcal{S}_{P_{1}P_{2}}$ as
@@
$\mathcal{H}_{u}=\ker(P_{1}P_{2}-I)=\mathcal{S}_{P_{1}P_{2}}.$
@@
First, we prove the following identity:
@@
This proves the inclusion in (3.3). We will now move on to the core part of the proof for the theorem. Since $\mathcal{H}_{cnu}=(\ker(P_{1}P_{2}-I))^{\perp}$, as a general fact, we have
@@
When discussing the product of projections, one immediately draws a connection with the classical von Neumann’s alternating orthogonal projections formula [20]. This formula relates to the limit of powers of products of two projections. We begin with recalling a notation. Given projections $P_{1}$ and $P_{2}$ defined on some Hilbert space $\mathcal{H}$, recall that the closed subspace $\mathcal{S}_{P_{1}P_{2}}$ is defined by (see (3.1))
@@
This proves the following, which appears to be new information to von Neumann’s projection formula:
@@
where $\mathcal{H}_{u}$ is the unitary part of the canonical decomposition of $P_{1}P_{2}$.
@@
then we say that $T$ is a $C_{\cdot 0}$ contraction. If both $T$ and $T^{*}$ are in $C_{\cdot 0}$, then we simply write [18, page 72]
@@ 5. Inner projections
This is also equivalent to the factorization of the orthogonal projection $P_{\mathcal{S}}$:
@@
for all $i\neq j$. This well-known identity will be used extensively in this paper.
@@
Clearly, the lcm function $\varphi$ is unique, and we simply express it as
@@
$\text{lcm}\mathcal{I}_{T}$ exists for all $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$.
@@
$\varphi\leq\psi\text{ whenever }\varphi H^{2}(\mathbb{D}^{n})\subseteq\psi H^{2}(\mathbb{D}^{n}),$
@@
This identity is true for all nonzero $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$. Consequently, if we apply this identity to the adjoint of $T$, which is $T^{*}$, then we see that
@@
Let $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ be a nonzero operator. Suppose $T$ is a product of two projections. By Theorem 5.3, we know that
$T=P_{\varphi_{T}H^{2}(\mathbb{D}^{n})}P_{\varphi_{T^{*}}H^{2}(\mathbb{D}^{n})}.$
Also, Theorem 3.1 yields
(5.4) $\ker(T-I)=H^{2}(\mathbb{D}^{n})_{u}=\varphi_{T}H^{2}(\mathbb{D}^{n})\cap\varphi_{T^{*}}H^{2}(\mathbb{D}^{n}).$
Corollary 5.4.
Proof.
$\{0\}=\varphi_{T}H^{2}(\mathbb{D}^{n})\cap\varphi_{T^{*}}H^{2}(\mathbb{D}^{n}),$
which is a clear contradiction as $\varphi_{T}\varphi_{T^{*}}\in H^{\infty}(\mathbb{D}^{n})$ is an inner function and
$\varphi_{T}\varphi_{T^{*}}\in\varphi_{T}H^{2}(\mathbb{D}^{n})\cap\varphi_{T^{*}}H^{2}(\mathbb{D}^{n}).$
This completes the proof of the corollary.
@@
This decomposition bears some resemblance to the structure of matrices that result from the finite products of projections (see Wu [23, Theorem 4.6] and Oikhberg [16]).

### v2

@@ Abstract.
We study operators that are products of two orthogonal projections. Our results complement some of the classical results of Crimmins and von Neumann. Particular emphasis has been given to projections associated with inner functions defined on the polydisc.
@@ 1. Introduction
for some projections $P_{1}$ and $P_{2}$ in $\mathcal{B}(\mathcal{H})$. Operators that admit the above factorizations have been examined in multiple contexts. Aronszajn, Browder, Dixnier, Kakutani, and Weiner are among the many more eminent names who have contributed to illuminate this subject (see [13, 22] for a thorough historical description). Two additional notable contributions relevant to our work are von Neumann’s formula for iterated products of projections and Crimmins’ analysis of products of two projections. Let us first revisit two characterizations from the list provided by Crimmins [5, 22] : An operator $T\in\mathcal{B}(\mathcal{H})$ is the product of two projections if and only if
@@
Given a closed subspace $\mathcal{S}$ of a Hilbert space $\mathcal{H}$, we denote the orthogonal projection of $\mathcal{H}$ onto $\mathcal{S}$ by $P_{\mathcal{S}}$. The preceding factorization of $T$ is referred to as the canonical factorization and will be crucial in the subsequent discussion. We present somewhat independent proofs of the above equivalence properties. Additionally, we offer the following new characterization: $T\in\mathcal{B}(\mathcal{H})$ is a product of two projections if and only if
@@
Before turning to von Neumann’s perspective, we pause for Nagy-Foias and Langer’s viewpoint on contractive operators acting on Hilbert spaces. This is relevant because, if $T\in\mathcal{B}(\mathcal{H})$ is a product of two projections, then $T$ is necessarily a contraction (that is, $\|Th\|\leq\|h\|$ for all $h\in\mathcal{H}$). Consequently, the theory of contractions [17]— an influential and impactful theory—applies to operators that are products of two projections. This paper presents some new perspectives on the structure of contractions that are products of two projections. Given a contraction $T\in\mathcal{B}(\mathcal{H})$, the celebrated canonical decompositions of $T$ is the orthogonal decomposition of closed subspaces
$\mathcal{H}=\mathcal{H}_{u}\oplus\mathcal{H}_{cnu},$
where $\mathcal{H}_{u}$ and $\mathcal{H}_{cnu}$ reduce $T$, and $T|_{\mathcal{H}_{u}}$ is unitary and $T|_{\mathcal{H}_{cnu}}$ is cnu (that is, $T|_{\mathcal{H}_{cnu}}$ on $\mathcal{H}_{cnu}$ does not have unitary summand). Moreover, the unitary summand $\mathcal{H}_{u}$ is given by
@@
This decomposition is due to Nagy-Foias and Langer [10, 18]. The spaces $\mathcal{H}_{u}$ and $\mathcal{H}_{cnu}$ are designated as the unitary part and the cnu part of $T$, respectively.
@@
respectively, where $\bigvee$ denotes the span closure of subspaces. The above results also complement the classical von Neumann’s alternating orthogonal projection formula. In fact, if $P_{1}$ and $P_{2}$ are projections, then the von Neumann’s alternating orthogonal projection formula tells us that
@@
where $\mathcal{H}_{u}$ is the unitary part of the canonical decomposition of the contraction $T=P_{1}P_{2}$. In the context of canonocial decomposition, we additionally include the following asymptotic properties of the cnu part of $T$ (see Theorem 4.2):
In view of the popular notations [17], we write the above asymptotic properties simply as:
@@
Now we look into a more concrete infinite dimensional Hilbert space, $H^{2}(\mathbb{D}^{n})$, the Hardy space of square summable analytic functions defined on the polydisc $\mathbb{D}^{n}$. The commutative Banach algebra of all bounded analytic functions on $\mathbb{D}^{n}$ is denoted by $H^{\infty}(\mathbb{D}^{n})$. Therefore, we have
@@
for all $f\in H^{2}(\mathbb{D}^{n})$. One knows that $T_{\varphi}\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ for all $\varphi\in H^{\infty}(\mathbb{D}^{n})$. The function $\varphi\in H^{\infty}(\mathbb{D}^{n})$ is said to be inner if $|\varphi|=1$ a.e. on $\mathbb{T}^{n}$ in the sense of radial limits. It is known that $\varphi$ is inner if and only if $T_{\varphi}$ is an isometry on $H^{2}(\mathbb{D}^{n})$. In this case, $T_{\varphi}T_{\varphi}^{*}$ is an orthogonal projection onto $\varphi H^{2}(\mathbb{D}^{n})$. In other words, we have
@@
(1.1) $P=P_{\varphi H^{2}(\mathbb{D}^{n})}.$
Here our goal is to classify operators that are products of two inner projections. Of course, our aim is to obtain an analytic answer to this problem. To achieve this, for each nonzero operator $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n})$, we associate an inner function that acts as the least common multiple within a suitable class of inner functions. Our approach is as follows: Given a bounded linear operator $T$ on $H^{2}(\mathbb{D}^{n})$, define the set (which appears to be a nonemtpy set)
@@
Under this notation, we also have
@@
In the spirit of canonical factorizations, Theorem 5.3 yields an analytic characterization of products of two inner projections: Let $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ be a nonzero operator. Then $T$ is a product of two inner projections if and only if
@@
Moreover, we observe in Corollary 3.3 that if $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ is a cnu contraction, then $T$ cannot be expressed as a product of two inner projections.
@@
We note that, unlike $\mathcal{I}_{T}$, the set $\mathcal{J}_{T}$ has the potential to be empty. Therefore, part of the above result also guarantees that the set $\mathcal{J}_{T}$ remains nonempty. In the final section of this paper, we illustrate some of our results through concrete examples as above.
The remaining part of the paper is arranged as follows: Section 2 primarily focuses on proving some of the well-known results, including those by Crimmins and Sebestyén. However, the more general framework we’ve provided here also emphasizes the possibilities for the larger problem. In Section 3, we explain the classical concept of canonical decompositions of contractions and relate them to our theory of the product of two projections. This provides a precise description of the orthogonal decompositions applicable to the operators under consideration. Section 4 establishes an additional connection between our theory and a classical result, providing new insights into von Neumann’s alternating orthogonal projection formula. In Section 5, we examine our underlying problem from a specific standpoint. We apply our theory to projections based on inner functions defined on the Hardy space over the polydisc. We continue to work on concrete situations in Section 6, where the projections are considered to be model projections. The final section, Section 7, illustrates some of our results on products of two projections using Blaschke products.
2. Crimmins’ perspective
In this section, we prove some results on operators that can be represented as products of two projections. As stated in the very beginning, this is a problem that has been triggered by many. We take Crimmins’ point of view and dedicate a part of the section to reproving his results. However, we address this within a marginally wider framework, anticipating more realistic uses in the structure of operators that can be expressed as products of operators.
@@ 2. Crimmins’ perspective
Let $T\in\mathcal{B}(\mathcal{H})$. If $T=PX$ for some projection $P\in\mathcal{B}(\mathcal{H})$ and linear operator $X\in\mathcal{B}(\mathcal{H})$, then
@@
Now, $T=P_{\overline{\text{ran}T}}T$, and hence $T=P_{\overline{\text{ran}T}}PX$. This implies $T=P_{\overline{\text{ran}T}}X$, which completes the proof of the lemma.
@@
Let $T\in\mathcal{B}(\mathcal{H})$. If $T=P_{1}XP_{2}$ for some projections $P_{1},P_{2}\in\mathcal{B}(\mathcal{H})$ and linear operator $X\in\mathcal{B}(\mathcal{H})$, then
@@
We are now presenting a result of Sebestyén [16, Corollary 2]. The present proof, compared to [16], is more elementary (the sufficient part) and is also presented for completeness. The result is clearly in line with Douglas’ range inclusion theorem.
@@
Crimmins once again receives credit for the equivalence of (1) and (4) [5, page 1595]. The present proof is once again different. In this context, we also refer the reader to [5, Theorem 3.1]. The same paper contains results (cf. [5, Corollary 3.8]) related to parametrizations and the uniqueness of pairs of projections whose products yield a given operator.
@@ 3. Nagy-Foias and Langer’s perspective
Note that if a bounded linear operator acting on a Hilbert space is a product of two projections, then it is necessarily a contraction. Consequently, the theory of contractions, widely recognized as being extremely rich, applies in this situation. This is a vast domain, and a thorough application of the theory of contractions to this particular class of operators and vice versa may not be entirely transparent. This requires a more in-depth investigation, which ought to be carried out. However, in this section, we will address a fundamental characteristic of the product of projections via the structure of contractions. To do that, we revisit the classic Nagy-Foias and Langer orthogonal decompositions of contractions [10, 18].
Let $T\in\mathcal{B}(\mathcal{H})$ be a contraction. Then there is a unique orthogonal decomposition $\mathcal{H}=\mathcal{H}_{u}\oplus\mathcal{H}_{cnu}$ such that both $\mathcal{H}_{u}$ and $\mathcal{H}_{cnu}$ reduce $T$, and $T|_{\mathcal{H}_{u}}$ is unitary whereas $T|_{\mathcal{H}_{cnu}}$ is completely non-unitary (cnu in short). This amounts to saying that $T|_{\mathcal{H}_{cnu}}$ does not have unitary summand (see Section 1). As a result, we get the diagonal decomposition:
@@
on $\mathcal{H}=\mathcal{H}_{u}\oplus\mathcal{H}_{cnu}$. Moreover, we have the representation of the unitary part $\mathcal{H}_{u}$ as [17, Chapt. 1, Theorem 3.2]
@@
This decomposition is popularly known as the Nagy-Foiaş and Langer or the canonical decomposition of contractions and stands as one of the most basic structures of contractions. In the following, we connect this with our theory of the product of two projections. For simplicity of notation, given projections $P_{1}$ and $P_{2}$, define the closed subspace $\mathcal{S}_{P_{1}P_{2}}$ as
@@
$\mathcal{H}_{u}=\ker(P_{1}P_{2}-I)=\mathcal{S}_{P_{1}P_{2}},$
where $\mathcal{H}_{u}$ is the unitary part of the contraction $P_{1}P_{2}$.
@@
First, we claim that
@@
This proves the inclusion in (3.3). We will now move on to the core part of the proof for the theorem. Since $\mathcal{H}_{cnu}=(\ker(T-I))^{\perp}$, as a general fact, we have
@@
We end this section with an application of Theorem 3.1 to a concrete situation, like inner projections defined on the polydisc (see the definition in (1.1)). Let $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$ be a nonzero operator. Suppose $T$ is a product of two inner projections, that is, there exist two inner functions $\varphi$ and $\psi$ in $H^{\infty}(\mathbb{D}^{n})$ such that
$T=P_{\varphi H^{2}(\mathbb{D}^{n})}P_{\psi H^{2}(\mathbb{D}^{n})}.$
Theorem 3.1 yields
(3.4) $\ker(T-I)=H^{2}(\mathbb{D}^{n})_{u}=\varphi H^{2}(\mathbb{D}^{n})\cap\psi H^{2}(\mathbb{D}^{n}).$
Corollary 3.3.
Proof.
$\{0\}=\varphi H^{2}(\mathbb{D}^{n})\cap\psi H^{2}(\mathbb{D}^{n}),$
which is a clear contradiction as $\varphi\psi\in H^{\infty}(\mathbb{D}^{n})$ is an inner function and
$\varphi\psi\in\varphi H^{2}(\mathbb{D}^{n})\cap\psi H^{2}(\mathbb{D}^{n}).$
This completes the proof of the corollary.
@@
When discussing products of two projections, one immediately draws a connection with the classical von Neumann’s alternating orthogonal projections formula [19]. This formula relates to the limit of powers of products of two projections. We begin with recalling a notation. Given projections $P_{1}$ and $P_{2}$ defined on some Hilbert space $\mathcal{H}$, the closed subspace $\mathcal{S}_{P_{1}P_{2}}$ is the common range space defined by (see (3.1))
@@
This proves the following, which appears to be new information to von Neumann’s alternating projection formula:
@@
where $\mathcal{H}_{u}$ is the unitary part of the canonical decomposition of the contraction $P_{1}P_{2}$.
@@
then we say that $T$ is a $C_{\cdot 0}$ contraction. If both $T$ and $T^{*}$ are in $C_{\cdot 0}$, then we simply write [17, page 72]
@@
$T=\begin{bmatrix}T_{u}&0\\ 0&T_{cnu}\end{bmatrix},$
where $T_{u}:=T|_{\mathcal{H}_{u}}$ is the unitary part of $T$. This immediately implies that
@@ 5. Inner projections
This is also equivalent to the factorization of the projection $P_{\mathcal{S}}$:
@@
for all $i\neq j$. This well-known identity will be used extensively in what follows.
@@
Clearly, the lcm function $\varphi$ is unique (if exists), and we simply express it as
@@
$\text{lcm}\mathcal{I}_{T}$ exists for all nonzero $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$.
@@
$\varphi\leq\psi\text{ whenever }\varphi H^{2}(\mathbb{D}^{n})\supseteq\psi H^{2}(\mathbb{D}^{n}),$
@@
This identity is true for all nonzero $T\in\mathcal{B}(H^{2}(\mathbb{D}^{n}))$. Consequently, if we apply this to $T^{*}$, then we see that
@@
This decomposition bears some resemblance to the structure of matrices that result from the finite products of projections (see Wu [22, Theorem 4.6] and Oikhberg [14]).